OpenStudy (anonymous):

M(-3,-1) is the midpoint of RS.If S has coordinates. (4, 6) what are the coordinates of R?

OpenStudy (studygurl14):

Midpoint formula: \[\frac{ x_1 + x_2 }{ 2 }, \frac{ y_1 + y_2 }{ 2 }\] Plug in what you know \[\frac{ 4 + x_2 }{ 2 } = -3\] \[\frac{ 6 + y_2 }{ 2 } = -1\] solve for x2 and y2

OpenStudy (anonymous):

Okay thanks

OpenStudy (studygurl14):

you're welcome